OpenStudy (anonymous):

A system of equations is shown below. 4x = - 3y + 17 3x - 4y = - 6 What is the solution to this system of equations? (-2, -3) (3, 2) (2, 3) (-3, -2)

OpenStudy (anonymous):

@kelliegirl33 ?

OpenStudy (anonymous):

4x = -3y + 17 ---> 4x + 3y = 17 now we have : 4x + 3y = 17 -->(-3) 3x - 4y = -6 --> (4) ------------------ -12x - 9y = - 51 -- result of multiplying by -3) 12x - 16y = - 24 -- result of multiplying by 4) -----------------add - 25y = - 75 y = -75/-25 = 3 now sub 3 in for y in either of the above equations 3x - 4y = -6 3x - 4(3) = -6 3x - 12 = -6 3x = -6 + 12 3x = 6 x = 6/3 = 2 x = 2 and y = 3 or (2,3)